# Logical computation demonstrated with a neutral atom quantum processor

### v2

@@ V Encoded Bernstein-Vazirani algorithm
VI Repeated loss correction and transversal operations for the $\llbracket 4,2,2\rrbracket$ code
VI.1 Repeated loss correction with CZ and dual CZ logical gates
@@
The circuits considered so far have only used loss correction and error detection at the very end, when decoding the final measurements. However, it’s important to be able to detect and correct errors between computational steps of an algorithm, to enable deeper computation. To that end, we explore using loss correction in the middle of a larger circuit, demonstrating repeated fault-tolerant loss correction. In order to make the circuits more interesting, between rounds of loss correction we apply some of the fault-tolerant logical operations given in Sec. III.
@@
Figure 12: The top circuit (a) measures the parity of four qubits, i.e., the syndrome of $ZZZZ$. However, it is not fault tolerant; a single $Z$ fault on the ancilla can spread undetected to a weight-two $Z$ error. The next circuit (b) remedies this, and simultaneously measures $XXXX$, using a trick from Rei (20). Now the bad $Z$ fault will be detected, or “flagged,” by the $XXXX$ measurement, and similarly the $ZZZZ$ measurement will flag a bad $X$ fault. This circuit gives an efficient, fault-tolerant error detection procedure for the $\llbracket 4,2,2\rrbracket$ code. However, it does not correct for a lost qubit. The bottom circuit (c) fixes this, halfway, by switching the direction of the final two CNOT gates. (This is equivalent to inserting swap-based “leakage-detection units” CBO+ (24).) The last two data qubits are replaced with fresh ancillas, so in particular if one of them is lost it will be replaced. Thus we can replace all four qubits in two rounds of error detection. Future work will investigate this paradigm by combining mid-circuit measurement, qubit reuse and reinitialization, along with periodic atom cooling to remove a fundamental obstacle to neutral atom quantum computing by limiting the depth of circuit any atom must undergo.
@@
Figure 13: (a) Encoded circuit that applies two rounds of encoded CZ gates to ${|\overline{++}\rangle}$, encoded in the $\llbracket 4,2,2\rrbracket$ code, with two rounds of error detection. Together, the two error detection rounds refresh all four data qubits. (b) The corresponding logical circuit. Since the CZ gates cancel out, both measurements should give $+$. (c) The baseline circuit that we run for comparison. The $X$ echo pulses cancel out coherent dephasing errors from the CZ gates and movement. Both measurements should give --. This baseline performs better than applying the two CZ gates sequentially without leaving the interaction zone.
We also ran a very similar experiment, except applying logical dual CZ gates, $(H\otimes H)\,\text{CZ}\,(H\otimes H)$, using $S_{X}\otimes S_{X}^{\dagger}\otimes S_{X}^{\dagger}\otimes S_{X}$.
@@
The encoded experiments can be compared against corresponding unencoded baselines. Given the system architecture, there are a few reasonable baselines from which to choose. For example, we may choose to return qubits to the storage zone between each CZ, or not. We have chosen to select unencoded baselines in which, as shown in Fig. 13(a), between each pair of CZs or dual CZs, both qubits are moved back to the storage zone and Pauli $X$ or $Z$ gates, respectively, are applied to each of them to cancel coherent dephasing from the movement and two-qubit gates. These baselines performed the best of those that we tried, disadvantaging any comparison between encoded and unencoded.
Results for both repeated logical CZ and repeated logical dual CZ, for up to 10 logical operations and 9 rounds of loss correction, are shown in Fig. 14, with detailed data in Table 6. Since the two physical baselines compile to the same circuit, we have combined them into a single baseline for comparison. The encoded repeated logical CZ experiments largely beat the corresponding baselines, including, for example, an encoded circuit consisting of 10 logical CZs with 9 loss-correction rounds outperforming an unencoded sequence of 10 CZs. Surprisingly, the repeated dual CZ experiment performs much worse for depths $8$ and $10$. Since it should be similar to repeated CZs, this could be from a mistake in scheduling the circuits, or possibly from machine parameter drift.
@@
Figure 14: Error rates for $\llbracket 4,2,2\rrbracket$ code repeated logical CZ and repeated logical dual CZ experiments, and the corresponding physical baseline, conditioned on no initial loss. Error detection between logical gates (Fig. 13) permits mid-circuit correction of atom loss.
Table 6: Data for our $\llbracket 4,2,2\rrbracket$ repeated controlled-$Z$ and repeated dual controlled-$Z$ experiments. The encoded experiment with CZ or dual CZ depth $k+1$ has $k$ loss-correction rounds. It uses $5+2k$ physical qubits, with $5+8k$ physical CZ gates. A trial is prerejected if the cat state verification fails, and it is postrejected if an uncorrectable error is detected. Results for all available data, and for data in which there is no initial loss are reported separately. The unencoded baseline experiments, the same for CZ as for dual CZ, are on two qubits and include echos to cancel coherent errors. Trials with initial loss are prerejected, and those with circuit loss are postrejected. Restricting to trials with no initial loss has no effect on the baseline.
      All data
CZ or dual Loss-correction
CZ depth rounds Trials Prerejected Postrejected $\Pr[\text{accept}]$ Errors Error rate
Encoded CZs 2 1 3400 189 732 $73_{-2}^{+1}\%$ 34 $1.4^{+0.5}_{-0.4}\%$
4 3 4600 587 2030 $43(1)\%$ 43 $2.2^{+0.7}_{-0.6}\%$
6 5 3700 290 2408 $27(1)\%$ 41 $4(1)\%$
8 7 3700 302 2674 $20(1)\%$ 37 $5^{+2}_{-1}\%$
10 9 3700 318 2979 $10.9(1)\%$ 21 $5(2)\%$
Encoded dual CZs 2 1 5120 633 1051 $67(1)\%$ 53 $1.5(0.4)\%$
4 3 8640 1388 3340 $45(1)\%$ 104 $2.7(0.5)\%$
6 5 5120 1156 2899 $21(1)\%$ 69 $6^{+2}_{-1}\%$
8 7 5120 1270 3544 $7.7^{+0.8}_{-0.7}\%$ 63 $16^{+4}_{-3}\%$
10 9 5120 1640 3242 $4.7(6)\%$ 38 $16^{+5}_{-4}\%$
Baseline 2 4960 293 302 $88.0(9)\%$ 77 $1.8(4)\%$
4 4960 200 431 $87.3(9)\%$ 113 $2.6_{-0.4}^{+0.5}\%$
6 4960 202 566 $84.5(10)\%$ 186 $4.4_{-0.6}^{+0.7}\%$
8 4660 202 654 $82(1)\%$ 210 $5.5_{-0.7}^{+0.8}\%$
10 4960 284 860 $77(1)\%$ 374 $9.8_{-0.9}^{+1.0}\%$
      Trials with no initial loss
CZ or dual Loss-correction
CZ depth rounds Trials Prerejected Postrejected $\Pr[\text{accept}]$ Errors Error rate
Encoded CZs 2 1 3119 127 603 $77_{-2}^{+1}\%$ 21 $0.9^{+0.4}_{-0.3}\%$
4 3 3787 197 1692 $50(2)\%$ 28 $1.5^{+0.6}_{-0.5}\%$
6 5 3209 205 2048 $30(2)\%$ 32 $3(1)\%$
8 7 2890 188 2026 $23(2)\%$ 29 $4^{+2}_{-1}\%$
10 9 3004 210 2417 $13(1)\%$ 15 $4(2)\%$
Encoded dual CZs 2 1 4282 333 746 $75(1)\%$ 31 $1.0^{+0.4}_{-0.3}\%$
4 3 6732 854 2324 $53(1)\%$ 58 $1.6^{+0.5}_{-0.4}\%$
6 5 4064 793 2302 $24(1)\%$ 48 $5(1)\%$
8 7 3626 775 2500 $9.7^{+1.0}_{-0.9}\%$ 50 $14^{+4}_{-3}\%$
10 9 3312 940 2171 $6.1(8)\%$ 26 $13^{+5}_{-4}\%$
VI.2 Random circuits
As explained in Sec. III, the $\llbracket 4,2,2\rrbracket$ code allows a variety of other fault-tolerant logical gates. Each of $\{\mathrm{CNOT},\mathrm{CZ},\mathrm{SWAP},H^{\otimes 2}\}$ can be implemented by permutation or transversal gates. This set generates a group of size $36$.
@@ VI.2 Random circuits
To test the performance of these gates, and in the spirit of Ref. Got (16), we design experiments by preparing encoded ${|00\rangle}$ and then applying up to four elements of the group at random, with error detection between each pair of elements. The sequences chosen are, using CNOT = CX for brevity:
@@
The encoding of sequence $j$ involves $j$ rounds of error detection. Note that the error detection circuits used here are those of Fig. 12(a), instead of those of Fig. 12(b); therefore they correct for lost qubits only at the end of the circuit. Future work may include executing these logical operation sequences with the circuits of Fig. 12(b).
The choice to error-detect between group elements is justified by imagining a computation with many [[4,2,2]] blocks. In that scenario, it is natural to write a circuit as alternating layers of CNOTs between blocks, and group elements within each block. It is then also natural to schedule error detection following each pair of layers in order to control the spread of errors among blocks.
The output of each random sequence is a stabilizer state with two generators. Each generator is measured to determine success of failure of the experiment. The encoded circuits can be directly compared to corresponding physical baselines that execute the same gate sequence. Results are shown in Fig. 15, with more details in Table 7. We find the encoded circuits consistently outperform the corresponding physical unencoded baselines.
@@
Table 7: Data for our $\llbracket 4,2,2\rrbracket$ random circuit experiments. The sequences of operations are specified in Eq. (1). The encoding of sequence $k$ includes $k$ rounds of error detection, $5+2k$ physical qubits and $5+8k$ physical CZ gates. For sequences 0, 1, and 2, the two stabilizer generators of the output each have weight one and are measured simultaneously in each experiment. An error is reported if either of the two generators have incorrect eigenvalue. The stabilizer generators of sequence 3 are $\{XZ,ZX\}$. The encoded circuit measures both of these generators in the same experiment via transversal $Y$-basis measurements. The physical baselines require separate experiments for each generator. Results for the two generators are therefore reported separately. All data is reported. Restricting to trials with no initial loss has no effect on the baseline.
    All data
Sequence Trials Pre. Post. $\Pr[\text{accept}]$ Errors Error rate
Encoded 0 1200 66 $42$ $91(2)\%$ 6 $0.6^{+0.6}_{-0.3}\%$
1 2800 119 $467$ $79_{-2}^{+1}\%$ 5 $0.2^{+0.4}_{-0.1}\%$
2 2500 132 $737$ $65(2)\%$ 9 $0.6^{+0.4}_{-0.3}\%$
$XZ$ 3 7200 595 $3292$ $46(1)\%$ 15 $0.5^{+0.3}_{-0.2}\%$
$ZX$ 3 7200 595 $3292$ $46(1)\%$ 24 $0.7(3)\%$
Baseline 0 1600 52 $51$ $94(1)\%$ 20 $1.3^{+0.7}_{-0.5}\%$
1 1600 40 $55$ $94(1)\%$ 13 $0.9^{+0.6}_{-0.4}\%$
2 1600 66 $47$ $93(1)\%$ 23 $1.6^{+0.7}_{-0.5}\%$
$XZ$ 3 1800 87 $108$ $89(1)\%$ 50 $3.1^{+0.9}_{-0.8}\%$
$ZX$ 3 1800 108 $134$ $87(2)\%$ 40 $2.6^{+0.9}_{-0.7}\%$
Trials with no initial loss
Sequence Trials Pre. Post. $\Pr[\text{accept}]$ Errors Error rate
Encoded 0 1104 45 $35$ $93_{-2}^{+1}\%$ 3 $0.3^{+0.5}_{-0.2}\%$
1 2645 95 $398$ $81_{-2}^{+1}\%$ 3 $0.1^{+0.2}_{-0.1}\%$
2 2289 105 $605$ $69(2)\%$ 9 $0.6^{+0.5}_{-0.3}\%$
$XZ$ 3 6315 425 $2702$ $50(1)\%$ 15 $0.5^{+0.3}_{-0.2}\%$
$ZX$ 3 6315 425 $2702$ $50(1)\%$ 24 $0.8(3)\%$
VII Repeated error correction with the distance-three Bacon-Shor code
Running deep logical circuits in a neutral atom quantum computer will ultimately require repeated correction of both errors within the qubit subspace and atom loss from erasure conversion. In order to demonstrate this, we choose the distance-three $\llbracket 9,1,3\rrbracket$ Bacon-Shor code. This code has previously been implemented on trapped-ion systems EDN+ (21); HBC (24). A code distance greater than two is needed to correct errors other than loss.
To our knowledge, the present work is the first demonstration of repeated error correction in neutral atoms. We employ a scheme similar to that of Steane and Knill Ste (97); Kni (05); PSR+ (24), leveraging its ability to correct atom loss; see Fig. 16. In this work, the ancilla qubits are held until readout at the end of the circuit, so each round of error correction requires the use of additional qubits. However, in principle, this scheme can be combined with mid-circuit measurement and reset where readout and loss detection, combined with mid-circuit reloading, will enable arbitrary-depth circuits.
[…]
After one round of error correction, we find a logical error rate of $4.9^{+1.0}_{-0.9}\%$ and after two rounds the error rate is $8^{+2}_{-1}\%$. Table 8 gives more details. While these results suggest the opportunity to further improve physical error rates, they also serve as a demonstration of the mechanism to correct for both errors and loss. This can be seen in the analysis of the successful trials with one round of error correction, where on average $0.20$ erasures and at least $0.14$ errors were corrected in the readout of the final data block, whereas $0.40$ erasures and $0.50$ errors were corrected in the previous round of $X$ and $Z$ error correction, a higher average number of corrections. In the successful trials with two rounds of error correction, on average $0.22$ erasures and at least $0.13$ errors were corrected in the readout of the final data block (consistent with the one-round data), while $0.95$ erasures and $0.88$ errors were corrected in the previous error-correction rounds.
In order to compare encoded and physical results, we must choose a meaningful physical baseline. From the standpoint of the encoded data, the logical circuit is a quantum memory, applying the logical identity operation. For a meaningful comparison, we show results for repeated teleportation of physical qubit states, following the pattern of repeated teleportation of the logical data in Fig. 16(a).
Table 8: Data for $\llbracket 9,1,3\rrbracket$ Bacon-Shor code repeated error correction (EC) experiments. The data for the encoded circuits includes all trials, including those with initial loss, and there is no preselection. When decoding logical $X$ ($Z$), a trial is postrejected if there are losses in two or more rows (columns) of the $3\times 3$ code, or if one row is lost and the other two disagree. For the unencoded baseline, a trial is prerejected if there are any initial losses and a surviving trial is postrejected if there are any circuit losses.
  Physical Physical
Experiment qubits CZs Trials Prerejected Postrejected $\Pr[\text{accept}]$ Errors Error rate
Encoded 1 EC round 27 36 2400 541 $77(2)\%$ 91 $4.9^{+1.0}_{-0.9}\%$
2 EC rounds 45 66 2400 1054 $56(2)\%$ 114 $8^{+2}_{-1}\%$
Baseline 2 1Q teleportations 3 2 $640$ 31 47 $88^{+2}_{-3}\%$ 7 $1.3^{+1.2}_{-0.7}\%$
4 1Q teleportations 5 4 $640$ 50 82 $79(3)\%$ 7 $1.4^{+1.3}_{-0.8}\%$
VIII Outlook
In this work, we have created large entangled logical states, implemented computational algorithms on encoded qubits, demonstrated qubit loss correction, and implemented a code with the capacity for repeated loss and error correction. Altogether, these results highlight the unique features and promise of neutral atoms as a platform for reliable fault-tolerant quantum computing.
The large encoded cat states and the Bernstein-Vazirani algorithms are enabled by the combination of scalability, high-fidelity operations inherent to neutral atoms, and all-to-all connectivity among the qubits. For larger-scale logical computation, this connectivity paired with high qubit count will enable the use of nonlocal error-correcting codes with high encoding rates and low physical qubit overhead XZZ+ (24).
All of the encoded experiments, particularly those with repeated error detection, highlight the usefulness of loss-to-erasure conversion in the context of encoded qubits. They demonstrate that logical circuits can operate effectively with atom loss, which will be a key feature of large-scale neutral atom quantum computers.
This work and similar work with other neutral atom platforms BEG+ (24), as well as recent work with superconducting qubits GC (24) and trapped ion qubits PSR+ (24); RAC+ (24), highlight a critical advance in the field of quantum computation: the transition from scientific results computed with physical qubits to computational advances achieved with logical qubits. After a relatively short time the neutral atom platform has shown significant advancement from its inception as an experimental quantum computing platform. In the near future we expect advances in both 2Q gate fidelities as more iterations are performed to reduce gate laser noise as well as significant scale up towards 10 0001000010\,00010 000 qubits allowing for larger distance codes, more performant logical qubits, and more impressive demonstrations of fault-tolerant, universal quantum computation.
The combination of a hardware-optimized qubit virtualization system, complete with high-rate and high-efficiency quantum error correction, with a programmable neutral atom quantum processor, inclusive of mid-circuit measurement NC+ (23), continuous atom reloading NKC+ (24), and increased two-qubit gate fidelity, is a promising path to achieving large-depth, logical computation for scientific quantum advantage.
Acknowledgments

### v3

@@ Appendix B Four-qubit codes
Appendix C Repeated loss correction and transversal operations for the $\llbracket 4,2,2\rrbracket$ code
C.1 Repeated loss correction with CZ logical gates
@@
The circuits considered so far have only used loss correction and error detection at the very end, when interpreting the final measurements. However, it is important to be able to detect and correct errors between the steps of an algorithm, to enable deeper computation. To that end, we demonstrate repeated fault-tolerant loss correction. In order to make the circuits more interesting, between rounds of loss correction we apply some of the fault-tolerant logical operations given in Sec. B.
Repeated error detection with distance-two codes has been demonstrated on ion traps, including three rounds on eight logical qubits SBA (24), and 15 rounds on four logical qubits YDKR (24). Demonstrations based on Gottesman’s protocol Got (16) for fault tolerance with the $\llbracket 4,2,2\rrbracket$ code, but without including repeated error detection, include WWJ+ (18); HF (19); Vui (18); LGL+ (17); CCNH (21). Repeated error correction, with higher-distance codes, has also been demonstrated on on trapped-ion processors RBL+ (21); PBP+ (24); PSR+ (24); RAC+ (24) and superconducting processors KLR+ (22); Goo (23); GC (24); PNH+ (24).
@@
Figure 13: a, Circuit to measure the parity of four qubits, i.e., the syndrome of $ZZZZ$. It is not fault tolerant; a single $Z$ fault on the ancilla can spread undetected to a weight-two $Z$ error. b, The next circuit remedies this, and simultaneously measures $XXXX$, using a trick from Rei (20). Now the bad $Z$ fault will be detected, or “flagged,” by the $XXXX$ measurement, and similarly the $ZZZZ$ measurement will flag a bad $X$ fault. This gives an efficient, fault-tolerant error detection procedure for the $\llbracket 4,2,2\rrbracket$ code. However, it does not correct for a lost qubit. c, The last circuit fixes this, halfway, by switching the direction of the final two CNOT gates MBG (23); CBO+ (24). The last two data qubits are replaced with fresh qubits, so in particular if one of them is lost it will be replaced. Thus all four qubits can be replaced in two rounds of error detection.
@@
Figure 14: (a) Circuit that applies two rounds of logical CZ gates to ${|\overline{++}\rangle}$, encoded in the $\llbracket 4,2,2\rrbracket$ code, with two rounds of error detection that refresh all four qubits. (b) The corresponding logical circuit. Since the CZ gates cancel out, both measurements should give $+$. (c) The baseline circuit run for comparison. The $X$ gates cancel out coherent dephasing errors from the CZ gates and movement. Both measurements should give --. This baseline performs better than applying the two CZ gates without leaving the interaction zone.
@@
The encoded experiment can be compared against a corresponding unencoded baseline. Given the system architecture, there are a few reasonable baselines from which to choose. For example, we may choose to return qubits to the storage zone between each CZ, or not. We have chosen to select a baseline in which, as shown in Fig. 14(a), between each pair of CZs, both qubits are moved back to the storage zone and Pauli $X$ gates are applied to each of them to cancel coherent dephasing from the movement and two-qubit gates. This baseline performed the best of those that we tried, disadvantaging any comparison between encoded and unencoded.
Results for up to 10 logical operations and 9 rounds of loss correction, are shown in Fig. 15. The encoded repeated logical CZ experiments largely beat the corresponding baselines, including, for example, an encoded circuit consisting of 10 logical CZs with 9 loss-correction rounds outperforming an unencoded sequence of 10 CZs.
@@
Figure 15: Error rates for $\llbracket 4,2,2\rrbracket$ code repeated logical CZ experiments, and the corresponding physical baseline, conditioned on no initial loss. Error detection between logical gates (Fig. 14) permits mid-circuit correction of atom loss.
C.2 Random circuits
@@
As explained in Sec. B, the $\llbracket 4,2,2\rrbracket$ code allows a variety of other fault-tolerant logical gates. Permutations and transversal gates can implement $\mathrm{CNOT},\mathrm{CZ},\mathrm{SWAP}$ and $H^{\otimes 2}$, generating a group of size $36$. To test the performance of these gates, and in the spirit of Ref. Got (16), we design experiments by preparing encoded ${|00\rangle}$ and then applying up to four elements of the group at random, with error detection between each pair of elements. The sequences chosen are, using CNOT = CX for brevity:
@@
The encoding of sequence $j$ involves $j$ rounds of error detection. Note that the error detection circuits used here are those of Fig. 13(a), instead of those of Fig. 13(b); therefore they correct for lost qubits only at the end.
The choice to error-detect between group elements is justified by imagining a computation with many $\llbracket 4,2,2\rrbracket$ blocks. In that scenario, it is natural to write a circuit as alternating layers of CNOTs between blocks, and group elements within each block. It is then also natural to schedule error detection following each pair of layers in order to control the spread of errors among blocks.
The output of each random sequence is a stabilizer state with two generators. Each generator is measured to determine success of failure of the experiment. The encoded circuits can be directly compared to corresponding physical baselines that execute the same gate sequence. As shown in Fig. 16, the encoded circuits outperform the corresponding physical baselines.
@@
Appendix D State preparation with the $\llbracket 16,6,4\rrbracket$ code
Figure 17: a, Fault-tolerant preparation of encoded $|{+}\rangle^{\otimes 3}$ in the $\llbracket 8,3,2\rrbracket$ code. The blue part uses the gadget in Fig. 13(b) to verify two of the stabilizers and swap qubits 3, 4 with 8, 9. The prepared state on qubits 0 to 7 is accepted only if qubits 8 and 9 are present and have trivial measurement outcomes. b, Non–fault-tolerant preparation. c, Layout of 250 atoms for preparing 25 copies of encoded $|{+}\rangle^{\otimes 3}$ using a. Colors only indicate the partition. Each $4\times 2$ tile joins a nearest same colored $2\times 1$ tile. d, Layout of 256 atoms for preparing 32 copies of encoded $|{+}\rangle^{\otimes 3}$ using b. e, Histogram of number of correct blocks of encoded $|{+}0{+}0{+}0\rangle$ in the $\llbracket 16,6,4\rrbracket$ code per shot, including shots with both $X$ and $Z$ basis measurements. Vertical lines indicate averages of 3.85(1) and 11.44(3) for the fault-tolerant and non–fault-tolerant protocol, respectively.
We use all the 256 atoms to prepare error-corrected logical qubits encoded in blocks of the $\llbracket 16,6,4\rrbracket$ tesseract code RAC+ (24). To this end, we first prepare multiple copies of $|{+}\rangle^{\otimes 3}$ encoded in the $\llbracket 8,3,2\rrbracket$ color code Cam (16), and then randomly pair them up, each pair being a copy of $|{+}0{+}0{+}0\rangle$ encoded in the tesseract code RAC+ (24).
We consider both fault-tolerant and non–fault-tolerant preparations of encoded $|{+}\rangle^{\otimes 3}$ in the $\llbracket 8,3,2\rrbracket$ code, as in Fig. 17(a) and Fig. 17(b), respectively. Compared to Fig. 17(b), Fig. 17(a) uses two extra qubits to measure the values of two stabilizers of the encoded $|{+}\rangle^{\otimes 3}$, and rejects upon any nontrivial measurement outcome or detected loss. For the non–fault-tolerant protocol, we first prepare 32 copies of encoded $|{+}\rangle^{\otimes 3}$, using the layout in Fig. 17(d), and then randomly pair them up into 16 copies of encoded $|{+}0{+}0{+}0\rangle$. For the fault-tolerant protocol, we instead prepare 25 copies of encoded $|{+}\rangle^{\otimes 3}$ (with six atoms unused), as shown in Fig. 17(c), and only pair up accepted ones.
In both protocols, we measure all the atoms of the prepared tesseract code blocks in the same $X$ or $Z$ basis, and then decode. Specifically, an encoded $|{+}0{+}0{+}0\rangle$ block can be either “rejected” due to $\geq 2$ errors or losses detected, or “correct” if all the three deterministic values of logical $X$ or $Z$ operators are trivial, or “wrong” otherwise.
For the fault-tolerant protocol, we run 1980 and 2000 trials for $X$ and $Z$ basis, respectively; for the non–fault-tolerant protocol, we run 960 and 1000 trials, respectively. For each of these trials, we repeat the random pairing of the (accepted) blocks of encoded $|+\rangle^{\otimes 3}$ independently for 100 shots. Results of these shots are shown in Fig. 17(e). Specifically, we calculate the error rate, defined as the ratio of the number of wrong blocks over the number of correct or wrong blocks, both numbers summing over all shots of $X$ and $Z$ bases. We observe an error rate of $0.53(1)\%$ for the fault-tolerant protocol, nearly one order of magnitude smaller than $3.98(3)\%$ for the non–fault-tolerant protocol. However, the average number of correct blocks per shot is 3.85(1) for the fault-tolerant protocol, compared to 11.44(3) for the non–fault-tolerant protocol. This demonstrates a tradeoff between error rate and acceptance rate enabled by loss conversion.